OpenStudy (thedolorosablu3capricousg1rl):

Stumped once again..) The distance from Carrie's house to the park was 3/4 of a mile. Carrie walked 2/3 of the way before she stopped for a drink. How far had Carrie walked? A. 5/7 mile B. 1/3 mile C. 1/2 mile d. 1/4 mile

OpenStudy (aveline):

What is 2/3 multiplied by 3/4?

OpenStudy (alekos):

well that would be two thirds of three quarters of a mile

OpenStudy (thedolorosablu3capricousg1rl):

well, if I were to multiply 2/3 and 3/4 wouldn't I need a common denominator? which would be 12??

OpenStudy (aveline):

When you multiply or divide, you do not need a common denominator. You only need a common denominator when you add or subtract.

OpenStudy (aveline):

\[\frac{ 2 }{ 3 }\times \frac{ 3 }{ 4 }\]

OpenStudy (thedolorosablu3capricousg1rl):

ok then, it would be 6/12, simplified to the lowest it can go would be 1/2

OpenStudy (aveline):

Yes!

OpenStudy (thedolorosablu3capricousg1rl):

thank you so much! I appreciate you helping me.